OpenStudy (anonymous):

(X-4)7^log7(x^2-3x-28)-log7(x^2-16)

OpenStudy (anonymous):

If you don't mind, can you retype your question? i dun understand (x-4)7^log...=7(x-4)^log...?

OpenStudy (anonymous):

(x-4)(x^2-3x-28) - log7(x-4)(x+4) =(x-4)(x-7)(x+4) - log7(x-4)-log7(x+4) That is all i know to do >.<